Eric’s average income for the first 4 months of the year is $1,450.25, what must be his
average income for the remaining 8 months so that his average income for the year is
1
Expert's answer
2020-06-08T18:54:18-0400

"\text{Average of the first four months}=1450.25\\[1 em]
\text{ The average }=\frac{\text {Sum of the values}}{ \text { The number of values }} \\[1 em] 
1450.25=\frac{\text {The sum of the first four months}}{ \text { The number of months }} \\[1 em] 
1450.25=\frac{\text {The sum of the first four months}}{ \text { 4 }} \\[1 em] 
\text{The sum of the first four months}=4 \times1450.25=5801\\[1 em] 
\text {The question did not mention the average year if we assumed it}=1,780.75\\[1 em] 
1,780.75=\frac{\text {The sum of the first four months+The sum of the following eight months}}{ \text { The number of months }} \\[1 em] 
1,780.75=\frac{\text {5801+The sum of the following eight months}}{ \text { 12 }} \\[1 em] 

 \text {5801+The sum of the following eight months}=12\times1,780.75=21,369 \\[1 em] 
 \text {The sum of the following eight months}=21,369 -5801= 15,568 \\[1 em] 
 \text {The average of the following eight months}=\frac{15,568}{8}=\fcolorbox{red}{aqua}{1946}"